OpenStudy (anonymous):

A business has $15,000 to spend on airline tickets to travel to a conference. It wants 27 of its employees to attend. The business wants to buy as many business-class seats as possible. The business-class seats cost $700. The economy-class seats cost $375. How many of each type of ticket should the business purchase? 11 business class and 16 economy class 12 business class and 15 economy class 15 business class and 12 economy class 16 business class and 11 economy class

OpenStudy (anonymous):

@pooja195

OpenStudy (eyust707):

We have two things that we want to know, but are not yet known; amount of business class seats and amount of economy class seats. Since there are two unknowns we need to write two equations.

OpenStudy (eyust707):

"It wants 27 of its employees to attend."

OpenStudy (anonymous):

is it c @eyust707

OpenStudy (eyust707):

yes

OpenStudy (eyust707):

did you calc out each one individually?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Given the equation 5x + y = 7, which equation below would cause a consistent-independent system? 10x + 2y = 14 -15x - 3y = -6 5x + y = -7 6x + 2y = 7

OpenStudy (anonymous):

is this d

OpenStudy (eyust707):

yes

OpenStudy (eyust707):

good job

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

How many solutions will this system of equations have? x - 2y = 24 3x - 6y = 72 No solution Infinite solutions One solution Two solutions

OpenStudy (anonymous):

either a o. b @pooja195 @eyust707

OpenStudy (eyust707):



OpenStudy (eyust707):